# Teaching Machine Learning Fundamentals with LEGO Robotics

### v1

@@ Abstract
This paper presents the web-based platform Machine Learning with Bricks and an accompanying two-day course designed to teach machine learning concepts to students aged 12 to 17 through programming-free robotics activities. Machine Learning with Bricks is an open source platform and combines interactive visualizations with LEGO robotics to teach three core algorithms: KNN, linear regression, and Q-learning. Students learn by collecting data, training models, and interacting with robots via a web-based interface. Pre- and post-surveys with 14 students demonstrate significant improvements in conceptual understanding of machine learning algorithms, positive shifts in AI perception, high platform usability, and increased motivation for continued learning. This work demonstrates that tangible, visualization-based approaches can make machine learning concepts accessible and engaging for young learners while maintaining technical depth. The platform is freely available at https://learning-and-dynamics.github.io/ml-with-bricks/, with video tutorials guiding students through the experiments at https://youtube.com/playlist?list=PLx1grFu4zAcwfKKJZ1Ux4LwRqaePCOA2J.
@@ I Introduction
AI education is crucial for enabling K-12 students to understand, use, and develop AI systems [7]. This imperative has sparked substantial research into effective pedagogical approaches and prompted countries worldwide to develop AI curricula [4].
Most existing work focuses on ethics or AI applications, while underlying algorithms remain inaccessible due to their complexity. These technical aspects are typically introduced only at the university level, with AI systems often treated as ‘black boxes’ in K-12 education [22]. Yet, research with adult users has shown that revealing the internal mechanisms of machine learning models can significantly improve understanding [13]. Moreover, while most courses use block-based programming rather than classical programming [10, 25], this still introduces cognitive load and shifts students’ attention toward implementation details rather than the underlying concepts.
@@
We address this gap by developing a web based platform Machine Learning with Bricks that integrates interactive data and algorithm visualizations with LEGO robotics, a widely available tool in schools that requires no additional purchase, to teach algorithms such as KNN, linear regression, and Q-learning to students aged 12 to 17. Using this platform, we conducted a two-day workshop with children to evaluate the effectiveness of our approach. Our key contributions are as follows:
@@
    Empirical evaluation showing improved conceptual understanding of machine learning concepts, positive shifts in AI perception, strong accessibility and visual appeal of the proposed platform, and increased motivation for continued learning.
@@ III Learning Platform
We developed Machine Learning with Bricks, a web-based platform that enables hands-on machine learning experiments using LEGO Education SPIKE Prime kits. The platform is designed for both individual use at home and instruction in the classroom. It establishes wireless connectivity with the robotic kit, enabling real-time visualization of student-collected data and providing interactive implementations of basic machine learning algorithms: KNN, linear regression and Q-learning. This section introduces the platform design, followed by detailed descriptions of three robots designed to illustrate the algorithms: the Fruit Detector for KNN, the Pitcher for linear regression, and the Crawler for Q-learning.
@@ 5 Can explain it well enough to teach it to a friend
The post-course survey repeated all pre-course questions to measure changes in knowledge and AI perceptions. Additional items evaluated course design, engagement, and impact using a standard 5-point Likert scale (1 = strongly disagree, 5 = strongly agree). An open-ended question collected additional feedback from participants.
@@ VI Results
We present findings from pre- and post-course surveys completed by 14 workshop participants. Results are organized into four subsections examining learning outcomes, AI perceptions, course design effectiveness, and motivation for future AI engagement.
@@ VI-A Learning outcomes
As shown in Table II, participants’ self-assessed knowledge increased across all topics. The largest gains were observed for questions concerning machine learning paradigms, their distinctions, and their applications (L2 and L3), both of which showed statistically significant improvements ($p<0.01$). Similarly, knowledge of specific algorithms and related concepts, including KNN, linear regression, reinforcement learning, and the exploration-exploitation dilemma (L5, L6, L7, L8), also increased, with all gains reaching statistical significance ($p<0.05$). Topics with pre-existing higher familiarity, such as general machine learning concepts (L1) and the importance of data in machine learning (L4), also showed increases in mean scores, although these improvements did not reach conventional statistical significance. Overall, these results suggest that the course was particularly effective in enhancing participants’ understanding of specific machine learning paradigms and algorithms, while also reinforcing general concepts.
TABLE II: Self-assessed knowledge on a 5-point Likert scale from Table I. Sample sizes vary due to missing responses. Statistical significance determined using Wilcoxon signed-rank tests [28]. * $p\leq 0.05$, ** $p\leq 0.01$
@@
L3. Where each of the machine learning paradigms can be applied** 14 1.64 (0.93) 14 3.07 (1.27) 0.001
L4. Data and its importance in machine learning 13 2.46 (1.27) 13 3.54 (1.45) 0.057
L5. KNN** 12 1.25 (0.87) 13 3.23 (1.54) 0.002
L6. Linear regression** 13 1.23 (0.60) 13 2.92 (1.50) 0.004
L7. Reinforcement learning* 13 1.85 (1.14) 14 3.14 (1.23) 0.031
L8. Exploration-exploitation dilemma* 13 2.08 (1.19) 13 3.23 (1.48) 0.029
VI-B Perception of AI
Table III illustrates how participants’ perceptions of AI shifted between the pre- and post-assessment. While general domain-related terms, as well as descriptions of applications, perceptions, and abstract concepts, remained largely consistent, the word “danger” disappeared from post-assessment responses. At the same time, participants began using specific machine learning terms, including “agent,” “rewards,” “data,” “linear regression,” and “supervised learning”, which implies that participants developed a more precise and technical comprehension of underlying machine learning concepts.
@@ VI-B Perception of AI
Abstract concepts future, science, technology, learning, progress, development, try out progress, science, future, technology (x2), learning
@@ VI-C Course design
As shown in Table IV, participants rated the course highly across all evaluated aspects, with mean scores ranging from 4.00 to 4.54 on a 5-point Likert scale. In particular, the clarity and understandability of presentations and examples (S5) and the enjoyment of hands-on activities (S7) received the highest ratings (both 4.54), indicating strong engagement and effective instruction. Other aspects, including the mathematical content (S1), the clarity and interest of data visualizations (S2–S3), and platform usability (S4), were also rated positively, suggesting that the course was both accessible and well-structured.
@@ VI-D Motivation for AI
In addition, the course appeared to increase participants’ interest in pursuing a career in AI. As shown in Fig. 8, the number of participants expressing interest in an AI/machine learning/robotics career rose from 5 pre-course to 8 post-course. Similarly, participants’ thematic interests shifted: in the post-survey, all participants indicated an interest in technology and science, compared to 11 out of 14 in the pre-survey.
@@ VII Discussion
This study evaluated the effectiveness of a web-based platform and two-day workshop curriculum to teach machine learning concepts to K-12 students through programming-free robotics activities. Our findings demonstrate that combining interactive visualizations with tangible LEGO robotics can successfully make basic machine learning algorithms accessible and engaging for young learners. This section interprets our key findings and acknowledges the limitations of the study.
@@ VII-A Effectiveness of Visualization-Based Robotics Approach
Significant improvements in students’ understanding of machine learning paradigms and specific algorithms (L2, L3, L5-L8) support our hypothesis that tangible, visualization-based approaches can effectively teach technical machine learning concepts to young learners. The particularly strong gains in understanding machine learning paradigms and their applications suggest that the hands-on, multi-algorithm approach helped students develop a broader conceptual framework for machine learning, rather than isolated knowledge of individual techniques.
@@ VII-B Shift to Technical Understanding
The emergence of technical machine learning terminology in post-course responses (agent, rewards, data, linear regression, supervised learning) indicates that participants developed a more precise conceptual understanding. The shift from general descriptors to algorithm-specific language suggests that students moved beyond abstract perceptions of AI toward concrete comprehension of its underlying mechanisms. This aligns with research showing that revealing the inner workings of machine learning models can improve understanding [13].
@@ VII-C Platform Design and Usability
The high scores for platform usability (D4: M = 4.00), visualization clarity (D2: M = 4.08), and instructional comprehensibility (D5: M = 4.54) support our design decisions to eliminate programming requirements and prioritize intuitive interfaces. Consistently positive feedback on hands-on activities (D7: M = 4.54) demonstrates that LEGO robotics successfully served as an engaging medium for abstract concepts, supporting the constructionist approach [19, 18].
@@ VII-D Motivation and Career Interest
The increase in AI career interest (from 5 to 8 participants) and universal post-course interest in technology and science suggest the workshop successfully sparked curiosity about machine learning and related fields. The high motivation scores (E1: M = 3.92, E2: M = 3.69) indicate that participants recognized both the intrinsic interest and practical value of machine learning skills, which may support sustained engagement beyond the workshop.
@@ VII-E Limitations
Despite these limitations, our findings provide initial evidence that the Machine Learning with Bricks platform and curriculum can effectively introduce machine learning concepts to young learners while fostering engagement and motivation for continued learning.
@@ VIII Conclusion
This paper presented Machine Learning with Bricks, an open-source web-based platform that combines interactive visualizations with LEGO robotics to teach core machine learning algorithms to K-12 students without requiring programming skills. Through a two-day workshop with 14 students aged 12-17, we demonstrated that this approach can effectively make basic machine learning concepts, such as KNN, linear regression, and Q-learning, accessible and engaging for young learners.
Our empirical evaluation revealed four key outcomes. First, participants achieved significant improvements in understanding machine learning paradigms and specific algorithms, with particularly strong gains in areas that were novel to them. Second, students’ perceptions of AI shifted from abstract apprehension toward concrete technical understanding. Third, the platform received high usability ratings and the hands-on activities were consistently engaging. Fourth, the workshop increased students’ motivation for continued AI learning and interest in pursuing AI-related careers.
These findings contribute to the growing body of research on AI education. By eliminating programming requirements and leveraging tangible robotics with real-time visualization, we show that young learners can grasp machine learning concepts and algorithms while maintaining high engagement and motivation.

### v2

@@ Abstract
This paper presents the web-based platform Machine Learning with Bricks and an accompanying two-day course designed to teach machine learning concepts to students aged 12 to 17 through programming-free robotics activities. Machine Learning with Bricks is an open source platform and combines interactive visualizations with LEGO robotics to teach three core algorithms: KNN, linear regression, and Q-learning. Students learn by collecting data, training models, and interacting with robots via a web-based interface. Pre- and post-surveys with 14 students indicate statistically significant improvements in self-reported understanding of machine learning algorithms, changes in AI-related terminology toward more technical language, high platform usability, and increased motivation for continued learning. This work suggests that tangible, visualization-based approaches can make machine learning concepts accessible and engaging for young learners while maintaining technical depth. The platform is freely available at https://learning-and-dynamics.github.io/ml-with-bricks/, with video tutorials guiding students through the experiments at https://youtube.com/playlist?list=PLx1grFu4zAcwfKKJZ1Ux4LwRqaePCOA2J.
Supplementary Materials
The survey instruments, anonymized response data, and analysis code used to evaluate the course and platform are publicly available at https://github.com/learning-and-dynamics/ml-with-bricks/tree/main/survey.
@@ I Introduction
AI education is crucial for enabling K-12 students to understand, use, and develop AI systems [7]. This imperative has sparked research into effective pedagogical approaches and prompted countries worldwide to develop AI curricula [4].
Most existing work focuses on ethics or AI applications, while underlying algorithms remain inaccessible due to their complexity. These technical aspects are typically introduced only at the university level, with AI systems treated as ‘black boxes’ in K-12 education [22]. Yet, research with adult users has shown that revealing the internal mechanisms of machine learning models can significantly improve understanding [13]. Moreover, while most courses use block-based programming rather than classical programming [10, 25], this still introduces cognitive load and shifts students’ attention toward implementation details rather than the underlying concepts.
@@
We address this gap by developing a web-based platform Machine Learning with Bricks that integrates interactive data and algorithm visualizations with LEGO robotics, a widely available tool in schools that requires no additional purchase, to teach algorithms such as KNN, linear regression, and Q-learning to students aged 12 to 17. Using this platform, we conducted a two-day workshop with children to evaluate the effectiveness of our approach. Our key contributions are as follows:
@@
    Empirical evaluation indicating statistically significant improvements in self-reported understanding of machine learning concepts, shifts in AI-related terminology toward more technical language, high platform usability, and increased motivation for continued learning.
@@ III Learning Platform
We developed Machine Learning with Bricks, a web-based platform that enables hands-on machine learning experiments using LEGO Education SPIKE Prime kits. The platform is designed for both individual use at home and instruction in the classroom. It establishes wireless connectivity with the robotic kit, enabling real-time visualization of student-collected data.
The platform provides interactive implementations of three basic machine learning algorithms: KNN, linear regression, and Q-learning. These were selected such that their underlying mathematical principles, such as coordinate systems, line equations, and tabular value updates, are accessible to students, allowing learners to develop mechanistic understanding rather than treating models as black boxes. Furthermore, the three algorithms represent distinct machine learning paradigms, providing broad coverage of foundational concepts.
This section introduces the platform design, followed by detailed descriptions of three robots designed to illustrate the algorithms: the Fruit Detector for KNN, the Pitcher for linear regression, and the Crawler for Q-learning.
@@ 5 Can explain it well enough to teach it to a friend
The post-course survey repeated all pre-course questions to measure changes in knowledge and AI-related terminology. Additional items evaluated course design, engagement, and impact using a standard 5-point Likert scale (1 = strongly disagree, 5 = strongly agree). An open-ended question collected additional feedback from participants.
@@ VI Results
We present findings from pre- and post-course surveys completed by 14 workshop participants. Results are organized into four subsections examining learning outcomes, AI-related terminology, course design effectiveness, and motivation for future AI engagement.
@@ VI-A Learning outcomes
As shown in Table II, participants’ self-assessed knowledge levels improved across all topics. The largest gains were observed for questions concerning machine learning paradigms, their distinctions, and their applications (L2 and L3), both of which showed statistically significant improvements ($p<0.01$). Similarly, self-assessed knowledge of specific algorithms and related concepts, including KNN, linear regression, reinforcement learning, and the exploration-exploitation dilemma (L5, L6, L7, L8), also increased, with all gains reaching statistical significance ($p<0.05$). Topics with pre-existing higher familiarity, such as general machine learning concepts (L1) and the importance of data in machine learning (L4), also showed increases in mean scores, although these improvements did not reach conventional statistical significance. Overall, these results suggest that the course was associated with gains in understanding of specific machine learning paradigms and algorithms, while also showing positive trends for more general concepts.
TABLE II: Self-assessed knowledge on a 5-point Likert scale from Table I. Sample sizes vary due to missing responses. Statistical significance determined using two-sided Mann-Whitney U tests. * $p\leq 0.05$, ** $p\leq 0.01$
@@
L3. Where each of the machine learning paradigms can be applied** 14 1.64 (0.93) 14 3.07 (1.27) 0.003
L4. Data and its importance in machine learning 13 2.46 (1.27) 13 3.54 (1.45) 0.063
L5. KNN** 12 1.25 (0.87) 13 3.23 (1.54) 0.001
L6. Linear regression** 13 1.23 (0.60) 13 2.92 (1.50) 0.003
L7. Reinforcement learning* 13 1.85 (1.14) 14 3.14 (1.23) 0.014
L8. Exploration-exploitation dilemma* 13 2.08 (1.19) 13 3.23 (1.48) 0.043
VI-B AI-Related Terminology
Table III illustrates changes in participants’ AI-related terminology between the pre- and post-assessment. While general domain-related terms, as well as descriptions of applications, perceptions, and abstract concepts, remained largely consistent, the word “danger” disappeared from post-assessment responses. At the same time, participants began using specific machine learning terms, including “agent,” “rewards,” “data,” “linear regression,” and “supervised learning”, which suggests increased familiarity with algorithm-specific vocabulary.
@@ VI-B AI-Related Terminology
Abstract concepts future, science, technology, learning, progress, development, try out progress, science, future, technology, tech, learning
@@ VI-C Course design
As shown in Table IV, participants rated the course highly across all evaluated aspects, with mean scores ranging from 4.00 to 4.54 on a 5-point Likert scale. In particular, the clarity and understandability of the presentations and examples (D5) and the enjoyment of hands-on activities (D7) received the highest ratings (both 4.54), suggesting strong engagement and positive reception of the instruction. Other aspects, including the mathematical content (D1), the clarity and interest of data visualizations (D2-D3), and platform usability (D4), were also rated positively, suggesting that the course was both accessible and well-structured.
@@ VI-D Motivation for AI
In addition, a higher number of participants expressed interest in pursuing a career in AI. As shown in Fig. 8, the number of participants expressing interest in an AI/machine learning/robotics career rose from 5 pre-course to 8 post-course. Similarly, participants’ thematic interests shifted: in the post-survey, all participants indicated an interest in technology and science, compared to 11 out of 14 in the pre-survey.
@@ VII Discussion
This study evaluated the effectiveness of a web-based platform and two-day workshop curriculum to teach machine learning concepts to K-12 students through programming-free robotics activities. Our findings suggest that combining interactive visualizations with tangible LEGO robotics can make basic machine learning algorithms accessible and engaging for young learners. This section interprets our key findings and acknowledges the limitations of the study.
@@ VII-A Effectiveness of Visualization-Based Robotics Approach
Statistically significant improvements in students’ self-assessed understanding of machine learning paradigms and specific algorithms (L2, L3, L5-L8) are consistent with our hypothesis that tangible, visualization-based approaches can effectively teach technical machine learning concepts to young learners. The particularly strong gains in understanding machine learning paradigms and their applications suggest that the hands-on, multi-algorithm approach helped students develop a broader conceptual framework for machine learning, rather than isolated knowledge of individual techniques.
@@ VII-B Shift to Technical Understanding
The emergence of algorithm-specific terminology in post-course responses (agent, rewards, data, linear regression, supervised learning) suggests increased familiarity with core machine learning concepts. The shift from general descriptors toward more technical language is consistent with research showing that revealing the inner workings of machine learning models can improve understanding [13], though more rigorous assessment would be needed to confirm deeper conceptual change.
@@ VII-C Platform Design and Usability
The high scores for platform usability (D4: M = 4.00), visualization clarity (D2: M = 4.08), and instructional comprehensibility (D5: M = 4.54) support our design decisions to eliminate programming requirements and prioritize intuitive interfaces. Consistently positive feedback on hands-on activities (D7: M = 4.54) suggests that LEGO robotics served as an engaging medium for abstract concepts, supporting the constructionist approach [19, 18].
@@ VII-D Motivation and Career Interest
The increase in AI career interest (from 5 to 8 participants) and universal post-course interest in technology and science suggest the workshop was associated with increased curiosity about machine learning and related fields. The high motivation scores (E1: M = 3.92, E2: M = 3.69) indicate that participants perceived both the intrinsic interest and practical value of machine learning skills, which may support sustained engagement beyond the workshop.
@@ VII-E Limitations
  • Lack of control group. Without a comparison condition, we cannot definitively attribute learning gains to our specific approach.
Despite these limitations, our findings provide initial evidence that the Machine Learning with Bricks platform and curriculum can serve as an engaging introduction to machine learning concepts for young learners, with promising indicators of motivation for continued learning.
@@ VIII Conclusion
This paper presents Machine Learning with Bricks, an open-source web-based platform that combines interactive visualizations with LEGO robotics to teach core machine learning algorithms to K-12 students without requiring programming skills. Through a two-day workshop with 14 students aged 12-17, we provide initial evidence that this approach can make basic machine learning concepts, such as KNN, linear regression, and Q-learning, accessible and engaging for young learners.
Our empirical evaluation revealed four key outcomes. First, participants achieved statistically significant improvements in self-assessed understanding of machine learning paradigms and specific algorithms, with particularly strong gains in areas that were novel to them. Second, students’ AI-related vocabulary shifted toward more algorithm-specific terminology. Third, the platform received high usability ratings and the hands-on activities were consistently engaging. Fourth, the workshop was associated with higher reported motivation for continued AI learning and interest in pursuing AI-related careers.
These findings contribute to the growing body of research on AI education. By eliminating programming requirements and leveraging tangible robotics with real-time visualization, we provide initial evidence that young learners can grasp machine learning concepts and algorithms while maintaining high engagement and motivation.